# Sensing single molecule magnets with nitrogen vacancy centers

### v1

@@ Acknowledgements
We thank O. Tal, J. S. Miller and L. Kronik for insightful discussions. We also thank I. Zohar and L. Schein-Lubomirsky for experimental assistance and proofreading the manuscript. AF acknowledges financial support from the Minerva Stiftung (Grant 714131) and the Israel Science Foundation (Grants 418/20, 419/20). AF also acknowledges support by the Kimmel Institute for Magnetic Resonance. AF is the incumbent of the Elaine Blond Career Development Chair in Perpetuity. JvS acknowledges funding from the German Science Foundation (DFG, SL104/10-1, SA1840/9-1). This research is made possible in part by the historic generosity of the Harold Perlman Family.

### v2

@@ Acknowledgements
We thank M. W. Doherty, O. Tal, J. S. Miller and L. Kronik for insightful discussions. We also thank I. Zohar and L. Schein-Lubomirsky for experimental assistance and proofreading the manuscript. IP is the incumbent of the Sharon Zuckerman research fellow chair. AF acknowledges financial support from the Minerva Stiftung (Grant 714131) and the Israel Science Foundation (Grants 418/20, 419/20). AF also acknowledges support by the Kimmel Institute for Magnetic Resonance. AF is the incumbent of the Elaine Blond Career Development Chair in Perpetuity. JvS acknowledges funding from the German Science Foundation (DFG, SL104/10-1, SA1840/9-1). This research is made possible in part by the historic generosity of the Harold Perlman Family.
Data Availability
The data that support the findings of this study are openly available at the following URL/DOI: 10.34933/400b149f-da31-441b-aace-0e4b037205b8
Supporting Information
The Supporting Information is available free of charge at the publisher’s website (TBD).
Contents: Simulation of NV effective sensing radius and volume (S1); Comparison between bulk and drop cast SMMs (S2); $T_{1}$ and $T_{2}$ curves fitting (S3); Additional statistics and control measurements (S4); Derivation of the coherence function in the frequency domain (S5); Model fitting to the data (S6); Temperature control at LT (S7); Sample preparation (S8); SMMs overview (S9); XPS analysis (S10); Experimental setup (S11); ab-initio calculations (S12); The T1 protocol (S13); Quantum sensing with NV centers (S14).